OpenStudy (anonymous):

Suppose that to make the golf team you need to score no more than 74 on average over 5 games. If you scored 77, 71, 77, and 67 in your first 4 games what is the highest score you can shoot in your 5th game and still make the team?

OpenStudy (anonymous):

let the last game score is x it says \[x+77+71+77+67\div5\le74\] solve for x..

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

how do you solve for x here

OpenStudy (anonymous):

ten is too large \[\frac{x+77+71+77+67}{5}\leq 74\] add the numbers in the numerator first

OpenStudy (anonymous):

ignore the "ten is too large" that came from some other post site is goofy these days

OpenStudy (anonymous):

\[\frac{x+292}{5}\leq 74\] multiply both sides by \(5\) then subtract 292